the sum of four consecutive even integers is -124. find the smallest.
Answer by mananth(16949)

let the even integers be x,x+2,x+4,x+6
their sum = x+x+2+x+4+x+6 =4x+12
4x+12 =-122
-12 -12
4x=-134
x=-33.5
..
-33.5,-31.5,-29.5-27.5
..
m.ananth@hotmail.ca